### v3

# JumpStarter: Getting Started on Personal Goals with Adaptive Personal Context Curation

Abstract.
Everyone aspires to achieve personal goals. However, getting started is often complex and daunting, especially for large projects. AI has the potential to create plans and help jumpstart progress, but it often lacks sufficient personal context to be useful. Personal context includes background information, preferences, and resources that can help tailor AI responses to individuals and their goals. We introduce JumpStarter, an AI-powered system that creates action plans and drafts by adaptively soliciting and integrating personal context. JumpStarter assists users by posing questions to elicit relevant context, breaking down goals into manageable steps based on context, and identifying relevant context to draft working solutions. A technical evaluation indicates that context curation results in plans and working solutions of higher quality. A user study demonstrates that compared to ChatGPT, JumpStarter significantly reduces mental load and increases efficiency for users. We discuss the design implications of adaptive personal context curation to better support users in complex problem solving.
context curation, planning, action initiation, personal goal management, productivity
ccs: Human-centered computing Interactive systems and tools
Figure 1. JumpStarter helps users get started on their personal goals through adaptive personal context curation. It first takes the user’s goal and elicits context for the goal. It then breaks down the goal into actionable subtasks. For each subtask, it helps users select relevant context and draft working solutions. It also aids users in refining these drafts by eliciting further context.
1. Introduction
Achieving personal goals is a common aspiration. Personal projects are part of everyday life, ranging from moving to a new apartment, to starting a YouTube channel, to applying for PhD programs. Getting started on these personal projects is crucial to turn intention into actual fulfillment (Sheeran and Webb, 2016; Gollwitzer, 1996). However, getting started on complex goals can be daunting and challenging. When faced with a new project, one needs to engage in sensemaking: constructing a mental model of the process and adapting it to their specific situation to figure out the actionable steps, all of which can be cognitively demanding (Johnson-Laird, 1983). Another major challenge is initiating concrete actions. For example, there is a significant gap between knowing when one needs to write a request email and actually being ready to write it (Sheeran and Webb, 2016). This shift requires concretization—such as defining the exact purpose of the request, deciding on the appropriate tone, and gathering the necessary information like relevant background details and supporting documents. Identifying and utilizing the right resources to operationalize the plan can be a significant hurdle to getting started.
Generative AI has the potential to create plans and help jumpstart progress, but it often lacks the personal context needed to be useful. Everyone has unique situations and requirements to consider when embarking on personal projects. For example, to get started on the goal of “applying for a PhD,” it is important to consider personal context, such as one’s academic background and geographic preferences. Moreover, a complex goal like this must be further decomposed to become actionable. It involves multiple tasks, such as researching universities and requesting recommendation letters, each of which requires further planning and coordination with other tasks. For example, requesting recommendation letters is itself a sub-goal that needs further breakdown—one must first identify recommenders, gather supporting documents, draft individualized request emails, send the emails, and follow up accordingly. Each of these tasks demands adaptive curation of relevant context to craft personalized solutions. For example, when drafting emails to recommenders, it is essential to solicit relevant context to ensure that the drafts are personalized to the recipient, which might include one’s relationship with the recommender, the programs being applied to, and aspirations for doctoral studies.
Popular AI tools like ChatGPT allow users to provide context throughout their interactions. However, it can be mentally demanding for users to recall and articulate the necessary context relevant to the task and to structure all the information within the bounds of prompts and chat history. Ideally, ChatGPT would remember all the context in the chat and automatically integrate the most relevant ones to support each task. However, this capability is limited by the inherent restrictions on context window size (Chen et al., 2023; Peng et al., 2023; Yen et al., 2024), and the context integration process is not transparent to users. Moreover, linear chat interfaces like ChatGPT lack the structure to help people systematically plan their roadmap toward their goals. Although people can ask ChatGPT to break down the steps further, it is easy to lose track of their progress and the context of each task.
Furthermore, theories about personal goals link their successful implementation to executive functions such as planning, task initiation, and working memory (Diamond, 2013), so approaches should be grounded in these functions. Planning involves a step-by-step breakdown of a broader goal (Kokkalis et al., 2013). This can be challenging as it requires identifying key tasks, estimating realistic timelines, and foreseeing potential obstacles, all of which demand significant effort (Teevan et al., 2016). While having a plan is crucial, it can sometimes remain abstract and lack the specificity and granularity required for tangible real-world actions (Gollwitzer, 1996; Kokkalis et al., 2013). Initiating concrete tasks demands heightened focus and meticulous attention to detail, which significantly increases cognitive load (Sweller, 2011). Effective execution of such tasks requires an active working memory, which aids in retaining and manipulating the relevant information needed to manage the task (Baddeley, 1992). Additionally, the Ovsiankina effect (Rickers-Ovsiankina, 1928) indicates that people are more likely to finish the task once they initiate it, underscoring the critical role of action initiation in goal completion.
To better support people in getting started on their personal goals, we introduce a system called JumpStarter (see Figure 1). Through mechanisms that adaptively solicit and integrate personal context, the system assists users by 1) decomposing tasks into hierarchical steps and 2) drafting working solutions. Here, the “working solution” refers to any tangible result that enables users to take real-world action. For example, a working solution could be a draft email to send to a recommender or a study schedule for the GRE. The process begins when a user inputs their personal goal. The system then poses questions to elicit the relevant context for the goal. Based on the user’s responses, the system breaks down the goal into subtasks the user can explore. If the system detects that a subtask is not actionable enough, it prompts the user, who can break it down further into additional subtasks. When the user is ready to work on a task, the system selects the relevant context and generates working solutions. If the user is not satisfied with the suggested solution, they can request more elicitation questions or provide their own prompts for refinement. Users can modify the suggested solution, and also save it as context to be used by the system in further subtasks. The user can then move to explore the next subtask, moving towards the completion of their goal.
This paper makes the following contributions:
  • JumpStarter, an AI-powered system for helping people get started on their personal goals with three main automated features: 1) adaptive personal context curation (including eliciting, saving, and selecting context), 2) hierarchical task decomposition, and 3) working solution generation.
  • A technical evaluation with and without context curation, showing the importance of context curation for getting plans and working solutions of high quality.
  • A user study showing that compared to ChatGPT, JumpStarter significantly reduces mental load and increases efficiency in getting started on personal projects.
We conclude by discussing the design implications of adaptive personal context curation to guide the design of future systems that support complex problem solving.
2. Related Work
2.1. Theory of Goal Achieving
Achieving personal goals is a complex process that involves various aspects of human psychology. This process encompasses numerous psychological factors, including motivation, self-belief, social support, feedback, planning, and more, as highlighted in theories like goal-setting theory (Locke and Latham, 2002) and the theory of planned behavior (Ajzen, 1991). To better understand how individuals navigate this intricate journey, action phase theory offers a comprehensive framework (Gollwitzer, 1990, 1996). It outlines the progression from initial desires to goal completion through distinct stages: pre-decisional, pre-actional (planning), actional, and post-actional phases, each critical in transforming aspirations into accomplishments. Action phase theory also emphasizes the importance of action initiation, as it marks the transition to the actional phase where individuals actively pursue their goals (Gollwitzer, 1990, 1996).
While other factors such as motivation and self-belief (Ajzen, 1991; Deci and Ryan, 2012; Maddux, 1995) are also crucial, we recognize the potential of computers in supporting the planning and action initiation aspects, especially for goals that primarily involve cognitive tasks or knowledge work (such as applying to PhD programs). In this work, we focus on providing support for individuals during the planning and action initiation phases of achieving their complex cognitive goals, helping them bridge the gap between mere goal intention and taking real-world action.
Planning presents its own set of challenges. In a probe study focusing on managing academic procrastination (Bhattacharjee et al., 2024), a majority of participants (13 out of 15) expressed a desire for structured action steps for their tasks. Many participants pointed out the struggles with unstructured or expansive tasks, indicating a need for planning support that includes step-by-step guidelines, structuring recommendations, and adaptive task breakdown. According to the implementation intention theory, plans that specify when, where, and how actions should be taken can transform abstract goals into concrete strategies, bridging the gap between intention and action (Gollwitzer, 1999). Similarly, goal-setting theory emphasizes the need for specific and actionable goals (Locke and Latham, 2002). Personalization is emphasized as another important dimension, suggesting that adapting goals to fit one’s real-life situations can significantly enhance the effectiveness of planning (Ekhtiar et al., 2023).
Action initiation is also challenging. A major hurdle is the need to take concrete actions—each task requires specific context to be acted on. However, people often juggle multiple tasks in their daily lives (Cypher, 1986). Frequent context switching between tasks can significantly drain cognitive resources, reducing the efficiency with which one can work on a project (Kersten and Murphy, 2006). To initiate action, context curation is necessary to create an enabling environment and to aid in concretization. The Fogg Behavior Model suggests that to persuade people to take action, one needs to provide facilitators (Fogg, 2009). An effective facilitator should inform people that the target behavior will not require a resource they do not have at that moment. Context curation—which involves creating a conducive environment tailored specifically for the task at hand—can thus facilitate the initiation of action (Weisberg et al., 2014).
2.2. Solutions to Support Action Planning
An action plan refers to a concrete, step-by-step breakdown of a task that helps people complete it  (Kokkalis et al., 2013). Prior studies show that people prefer not to make action plans for their own tasks (Krieger et al., 2009; Teevan et al., 2016), as the costs are immediate but benefits are deferred (Kaur et al., 2018). Therefore, it will be beneficial to provide people with support for action planning.
There are many HCI works focusing on community or crowd-based approaches to help people do a step-by-step breakdown of their goals. For example, TaskGenies (Kokkalis et al., 2013) decomposes tasks into concrete steps by using online crowds to create new plans and using algorithms to identify and reuse existing ones. PlanHelper (Liu et al., 2022) helps users construct activity plans (such as for bodybuilding or sightseeing) by processing answer posts in community-based QA platforms. Specifically for writing tasks, a vocabulary (Kaur et al., 2018) is introduced as a cognitive scaffold to enable other people who have the necessary context (such as collaborators), to create action plans for writers. In a collaborative teamwork setting, MixTAPE (Rahman et al., 2020) helps generate action plans for design teams by converting call notes and client briefs into prioritized, assigned tasks using a synthesis algorithm and machine learning models. While not directly supporting action planning, many other works, particularly in the crowdsourced field, focus on decomposing a large task into manageable micro-tasks that help people make progress with limited time and resources (Teevan et al., 2016; Cai et al., 2014; Williams et al., 2019).
With the advancement of generative AI techniques, large language models (LLMs) like GPT-4 (Achiam et al., 2023) hold the potential to provide intelligent support for planning. There is a series of NLP works that study the potential of automating real-world planning (Lal et al., 2024; Xie et al., 2024; Zheng et al., 2024). For example, a multi-agent approach (Lal et al., 2024) is introduced to determine the steps needed to complete a task (such as how to plant a garden), emphasizing the importance of tailoring the steps to meet users’ specific needs (such as planting a garden without using pesticides). However, studies found that the current language agents are not yet capable of automatically handling complex planning tasks (Xie et al., 2024; Zheng et al., 2024); humans still need to be involved to make meaningful plans for real-world tasks.
A popular online tool called Goblin Tools MagicTodo [^1] https://goblin.tools/ uses AI techniques to support people in breaking down their tasks layer by layer. Throughout our experiments with MagicTodo, we identified significant limitations, particularly in determining the appropriate level of task breakdown granularity. In theory, tasks can be endlessly decomposed into trivial steps, such as opening a browser. The challenge is to find a manageable granularity that remains meaningful for users. To address this, JumpStarter examines subtask detection to help users decide whether a task requires further breakdown while still allowing for personal control. A recent HCI work, ExploreLLM (Ma et al., 2024), also explores helping users decompose complex tasks into subtasks with LLMs. Their study shows that a schema-like structure is useful for people to do planning tasks, with ChatGPT as a baseline. The system also provides a dedicated input box to prompt users to enter their personal context and preferences for personalization. However, ExploreLLM only examines one layer of task decomposition and relies on users to provide context themselves. In JumpStarter, we explore multi-layer task decomposition and adaptive personal context curation throughout the process.
Figure 2. A screenshot of JumpStarter creating plans and drafting working solutions for the goal Apply for a PhD in NLP. (A) Task breakdown is shown as a subtask tree, with the goal being the root node. Subtasks decomposed from the same parent node are shown on the same level. (B) Saving the working solution draft. (C) Detailed descriptions of the selected subtask are shown. (D) The working solution is generated, considering the specification from the user – ”I want schools in midwest of US”. Users have three options to improve the draft: regenerate, add context and regenerate, and iterate based on users’ new specifications.
2.3. Solutions to Support Action with Context Curation
Context curation is a fundamental design concept that impacts productivity. It involves eliciting, saving, and selecting relevant information to tailor solutions for specific tasks and individuals (Kerne et al., 2014). The idea of ”readiness-to-hand” (Heidegger, 2010) illustrates how we perceive tools as extensions of our abilities, emphasizing the close intertwining of cognition and surroundings. For example, mise-en-place (Weisberg et al., 2014), a strategic arrangement of tools and ingredients before cooking, demonstrates how well-curated contexts can enhance efficiency. According to cognitive theories, our thoughts and actions are adaptive responses to our environment, underscoring the deep connection between human cognition and context (Winograd, 1986). This structural coupling highlights the significance of context curation in improving productivity and interactions with technology.
Context elicitation has been studied in the NLP field. Research in proactive information-seeking dialogue (Deng et al., 2023; Zhang et al., 2023) focuses on approaches to asking clarifying questions (Zhang et al., 2024; Malaviya et al., 2024) and eliciting user preferences (Zhang et al., 2018; Handa et al., 2024; Li et al., 2023) to better meet users’ information needs. However, this body of research often centers on pure information-seeking tasks, such as web searches or online purchases, and the experiments are often conducted in simulated chat environments. JumpStarter broadens this scope by facilitating interactions with real users across a wider variety of real-world tasks in a structured manner. Moreover, it advances this research by eliciting context beyond typical question-and-answer exchanges and incorporating personal files into the process, making context elicitation more practical.
Context saving and selection were studied in task-centric information management systems. In terms of context saving, UMEA (Kaptelinin, 2003) helps users create and manage project-specific work context by providing dedicated project spaces and monitoring user activities within a desktop environment. The Personal Project Planner (Jones et al., 2008) introduces an in-context creation feature that enables users to produce documents and emails within a project plan, automatically inserting hyperlinks to provide context to the task. Similarly, Towel (Conley and Carpenter, 2007) allows users to drag any semantic object (e.g., project, email, person) onto the to-do list, saving them as resources under the to-do item. Context selection has been particularly examined in programming scenarios. For example, Mylar (Kersten and Murphy, 2006) introduces a task context model to help programmers find and identify information relevant to their current task by utilizing task interaction history. JumpStarter not only streamlines the integration of context into users’ tasks but also dynamically optimizes task execution, transforming static context management into a proactive and adaptive workflow catalyst.
3. JumpStarter System
JumpStarter (see Figure 2) is an interactive system that helps users get started on their personal projects with adaptive personal context curation. The system takes the user-specified goal as the input, and outputs personalized plans and working solutions for the user. Here, a working solution refers to any tangible result that can assist users in taking real-world actions related to a task. For instance, a working solution for requesting recommendation letters might be an email draft to send to the advisor. Similarly, for passing the GRE test, a working solution could include a study schedule.
Overall, JumpStarter utilizes a hierarchical tree structure. We chose this approach to organize tasks because it offers a clear, logical flow that naturally mirrors the breakdown of complex goals into manageable components. This structure facilitates easy tracking of task dependencies and progression. Features such as collapsing or expanding nodes allow for customizable views, ranging from simple outlines to detailed oversight. Additionally, the ability to attach documents directly to task nodes enhances organization and accessibility, making this approach particularly effective for managing personal projects.
JumpStarter is powered by GPT-4222We used gpt-4-turbo throughout the system. and implemented with a Flask/Python web framework. In this section, we describe how our system works with an example walk-through, and the implementation details of the system.
3.1. System Description
As an example, consider John, who aims to apply for a PhD in NLP. In this subsection, we will demonstrate the typical interactions between John and our system through a detailed walk-through.
First, John establishes his goal by entering “Apply for a PhD in NLP” into a text box. He then clicks the Start button to submit it.
3.1.1. Context Elicitation for Goal
Based on John’s goal, the system uses the LLM to generate questions to elicit context from him (see Figure 3). For John, the system generates three questions: ”Can you provide your CV or Resume?”, ”Which universities or programs are you considering for your PhD in NLP?”, and ”Do you have letters of recommendation, or do you need guidance on how to obtain them?” Since John does not know which universities he might be applying to, and he has not obtained the recommendation letters, he only uploads his CV to the first question. Then, he clicks the Let’s get started button to proceed. This first step aims to clarify John’s goal by eliciting the necessary context that will be used globally throughout John’s interaction with JumpStarter, ensuring that the plans and working solutions are created based on John’s situations.
Figure 3. JumpStarter generates questions to elicit context from users to clarify the goal. The user uploads his CV.
3.1.2. Subtask Suggestion
Leveraging the context obtained from the first step, JumpStarter generates the initial subtask tree for John, as depicted in Figure 4E, which consists of a list of subtasks. It also presents John with an overview of the titles, descriptions, and duration of completion of these subtasks (see Figure 4F). After reviewing them, John gains brief insights into each subtask and decides to explore them in order.
Figure 4. The initial subtask tree and the overview for the goal Apply for a PhD in NLP. (E) The task breakdown for the goal. (F) The overview of all the subtasks of the goal, including the titles, descriptions, and duration of completion of the subtasks.
Figure 5. JumpStarter suggests further decomposition for the first subtask Identify Potential PhD Programs. (1) John presses the button Decompose the task. (2) JumpStarter generates the subtasks for John.
3.1.3. Subtask Detection
After reviewing the initial subtask tree, John thinks the list generally makes sense and proceeds to click on the first node: Identify Potential PhD Programs. As shown in Figure 5, the node is highlighted, and its title and description are shown on the top right of the interface. At the same time, the system shows a message as it detects this task needs to be further decomposed to make it actionable. The system offers two options for the user to select: the recommended option Decompose the task and the alternative option Directly start drafting. Clicking on the first button will further break down the selected task into several subtasks, while clicking on the second will generate a working solution for the task. Following the recommendation, John clicks the button Decompose the task, after which the subtasks for Identify Potential PhD Programs are generated, as shown in Figure 5.
3.1.4. Working Solution Drafting
John reviews all subtasks of Identify Potential Programs and proceeds to click on the first subtask Research Universities and Programs. JumpStarter deems this subtask actionable and suggests John directly start drafting the working solution for it. Following the suggestion, he obtained the initial working solution that contains a list of universities offering PhD programs in NLP. Upon reviewing the list, John realizes it includes many global institutions. However, he prefers schools in the midwest of the United States. To iterate, he enters ”I want schools in midwest of US” (see Figure 2D). After reading it, John is happy to have this list as the first draft of his school application list. Note that John is provided with three alternatives to further improve the draft, as shown in Figure 2D: 1) regenerate; 2) add more context and regenerate; and 3) specify details and iterate on the current draft. Subsequently, John copies and pastes the draft into the working solution section of the task Research Universities and Programs (see Figure 2B) and saves it, which appears as a file icon on the tree (see Figure 2A) with the subtask crossed out.
(a) Selecting relevant context for forking.
(b) Entity-based task decomposition (forking).
Figure 6. Context selection for forking and the task decomposition after applying forking on Identify Faculty Members.
3.1.5. Forking Detection and Context Selection
After obtaining the initial working solution for Research Universities and Programs, John moves on to the next subtask titled Identify Faculty Members. JumpStarter suggests that the subtask lacks the specificity needed for direct action and suggests further breakdown. John agrees and clicks the Decompose the task button. Following this, JumpStarter determines that the subtask should be forked to address the specifics of each program identified earlier. It then selects and presents the relevant context, specifically the working solution of the previously completed step Research Universities and Programs that includes a list of universities (see Figure 6(a)). John agrees with the suggestion and clicks the Accept button to proceed with forking using the selected context. Note that if John disagrees with the suggested context, he can select other context variables from a dropdown button; alternatively, he can also Decline and proceed with breakdown if he thinks that the subtask should not be forked. The updated subtask tree after forking is displayed in Figure 6(b). With the new subtasks of evaluating different programs, John proceeds to explore each program in greater detail by clicking on them sequentially.
3.1.6. Context Selection for Working Solution Generation
Next, John decides to explore the task Get Recommendation Letters and decomposes it as the system suggests. John selects the first subtask Compile a list of recommenders. JumpStarter uses his CV uploaded in the first step to generate a draft of recommenders: Prof. Blake White, Prof. Julian Deng, and Dr. Alice Feng. John likes the draft and saves it. Subsequently, John moves on to the subtask Reach out to potential recommenders, and JumpStarter breaks down the task based on the recommenders identified in the previous step. John then selects the subtask Reach Out to Potential Recommenders: Prof. Blake White, which requires “Send a personalized email to Prof. Blake White about your PhD aspirations.“ In addition to John’s CV, JumpStarter suggests two more relevant context: the information of each recommender (contains the relationship between Prof. Blake White and John) and the list of potential schools, as shown in Figure 7(a). Following the suggestion, John receives the initial email draft.
(a) Context selection for working solution generation.
(b) Context elicitation for working solution iteration.
Figure 7. Context selection and elicitation for drafting and iterating the working solution of the subtask Reach Out to Potential Recommenders: Prof. Blake White.
3.1.7. Context Elicitation for Working Solution Iteration
After reviewing the initial draft of the email for Prof. Blake White, John decides it needs improvement, though he is unsure how to proceed. He clicks the ”Add Context and Regenerate” button, prompting JumpStarter to generate several questions designed to draw more context from him (See Figure 7(b)). John then includes details about specific projects and papers he collaborated on with Prof. White and regenerates a refined draft. John likes the new version because it highlights specific aspects of his past collaborations with Prof. White, giving it a more personalized touch.
With all of these functionalities of JumpStarter, John manages to generate a plan and working solutions that he is satisfied with. This enables him to progress further and ultimately achieve his goal of applying for a PhD in NLP.
3.2. Data Representation
In our system, each context is represented as a key-value pair, where the key denotes the context’s name, and the value is its content. For instance, the pair “Location preference: Midwest of US” indicates the user’s preference for a specific location—Midwest of US.
There are two types of context in our system: global context and local context. Global context refers to context that is universally applied by default throughout the system. Specifically, it consists of all context elicited from the questions the system asks at the very beginning after the user enters their goal. The local context includes all saved working solution drafts and the context added by the user throughout their interaction with the system.
Additionally, we represent lists of subtask nodes generated by our system in the form of a tree, with the user’s goal serving as the root node. Each subtask node stores task-specific attributes, including working solution drafts, task titles, task descriptions, and estimated duration of completion, among others.
3.3. Feature Implementation
In this section, we provide implementation details for the key features of our system, as introduced in Section 3.1. Note that the user’s goal and the global context are integrated into the system prompt, which we use to implement all of our key features.
3.3.1. Context Elicitation for Goal
Based on the user’s input goal, JumpStarter poses questions to elicit context relevant to that goal from the user. To achieve this, we instruct GPT-4 to assess which existing documents the user might possess that could serve as the context. GPT-4 will suggest what file to upload along with an elicitation question. Our system then lists that question in the UI and allows the user to upload their file. Alternatively, if GPT-4 generates a question that does not require any documents from the user, an input text box will be provided in the UI, allowing users to directly type their responses. See the GPT prompts we used in Appendix A.1.
3.3.2. Subtask Suggestion
Given a task that the user decides to break down, JumpStarter decomposes it into a list of subtasks and attaches them as child nodes to the task node. We instruct GPT-4 to generate the list, where each subtask should have a name, associated descriptions, and an estimated duration for completion. This helps users understand the scope and duration of each task. To minimize the occurrence of repetitive subtask nodes across the tree, we prompt GPT-4 to consider the overall tree structure and suggest subtasks that naturally complement the existing hierarchy. See the GPT prompts we used in Appendix A.4.
3.3.3. Subtask Detection
Given a task that the user selects, JumpStarter evaluates whether it is actionable and sufficiently detailed. If it is not actionable, JumpStarter recommends further breakdown. If it is actionable, JumpStarter recommends that the user directly drafts the working solution. We use GPT-4 to perform subtask detection. To determine the optimal prompting method, we experimented with different prompting techniques. To optimize both the accuracy and user experience in terms of loading time, we decided to adopt Chain of Thought (CoT) prompting with few-shot examples and integrate the tree level of the target task into the prompt. More details about the experiments and results can be found in Section 4.1. See the GPT prompts we used in Appendix B.
3.3.4. Working Solution Drafting
Given a task that the user decides to work on, JumpStarter helps users draft working solutions, which contains the following parts.
Context Selection for Working Solution Generation
JumpStarter starts by selecting the relevant context for the given task. Specifically, given the title and detailed descriptions of the task, we prompt GPT-4 to select the most relevant context from the current local context collection. The selected context keys are presented as a checklist in a pop-up window in the system. Users are allowed to de-select any context key if necessary. Users can also add any context they deem useful from a drop-down list. We adopt the CoT prompting method to achieve this feature. See the prompt we used in Appendix A.3.1.
Working Solution Generation
The selected context is then applied to generate the working solution draft. We directly prompt GPT-4 to help with the task in a zero-shot manner by providing the task title, its descriptions and selected context as the input (along with the user’s goal and the global context in the system prompt). See the prompt we used in Appendix A.7.
Context Elicitation for Working Solution Iteration
As mentioned in Section  3.1, users can click on the “Add context and regenerate” button if they attempt to iterate on a working solution draft but are not sure how. The system will then perform context elicitation for working solution iteration for the user. Similar to context elicitation for goal (see Section 3.3.1), the system provides questions to elicit more context from users, hopefully enabling them to get a better draft. To achieve this, we prompt GPT-4 in the CoT manner. See the prompt we used in Appendix A.2. With the newly elicited context, the system then triggers the “working solution generation” feature to generate an iterated draft for users.
3.3.5. Forking Detection and Context Selection
Forking Detection
Given a task that requires further breakdown (as determined by the “Subtask Detection” feature), JumpStarter will then assess whether it requires forking, i.e., whether it needs to be decomposed based on different entities. Unlike the typical task breakdown that treats subtasks as sequential, forking focuses on decomposing subtasks in a way that allows them to be completed in parallel, independent of the order in which they are generated. If it is not forking, JumpStarter directly goes into a typical task breakdown without any changes in the UI. If it is forking, JumpStarter then triggers the context selection for forking feature (see the paragraph below). We prompt GPT-4 in a CoT fashion with four in-context examples. See the GPT prompts we used in Appendix A.6.
Context Selection for Forking
Given a task that requires forking (as determined by the “Forking Detection” feature), we instruct GPT-4 to select the most relevant context keys for forking. See the prompt we used in Appendix A.3.2.
4. Technical Evaluation
JumpStarter helps people create a hierarchical action plan and draft working solutions by detecting subtasks and curating personal context. In this section, we describe how we evaluated the accuracy of subtask detection and the impact of context curation on generated plans and working solutions.
4.1. Subtask Detection Experiments
4.1.1. Experiment settings
Subtask detection involves determining whether a task needs further decomposition to become actionable. With vanilla zero-shot and few-shot prompting as our baselines, we experimented with three methods for detecting subtasks with GPT-4 (see Appendix B for detailed prompts). We first experimented with few-shot prompting in the Chain-of-Thought (“CoT”) style (Wei et al., 2022). We then integrated the tree levels of the task nodes into the prompt, hypothesizing that nodes at shallower levels (e.g., the first level) would be less likely to be actionable than those at deeper levels (“Tree”). We also explored incorporating the initial solution draft into the prompt to encourage the system to suggest task decomposition depending on the quality of the draft (“Draft”). We evaluated the accuracy of these methods on an expert-labeled test suite. In all experiments, the temperature was set to $1$, the max tokens were set to $2048$, and top-P was set to $1$. We report the average across five runs for each experimental setting.
4.1.2. Test suite construction
Drawing inspiration from (Little, 1983), we created a test suite comprising four real-world task scenarios, each representing a distinct aspect of everyday life: 1) Applying for a PhD program (Academic), 2) Obtaining a driver’s license (Practical), 3) Finding a surfing camp (Recreational), and 4) Arranging a trip abroad (Travel). For each scenario, we created five distinct test cases, resulting in a total of 20 test cases. An example test case is shown below:
Compile a List of Potential Universities: start by identifying the universities that offer PhD programs in Natural Language Processing (NLP). Research and compile a comprehensive list of these universities.
We recruited four expert participants—one for each task scenario—via a university mailing list (average age=26.3, two female, two male). These participants reported having completed the tasks in the past six months and were compensated \$10 for their participation. For each test case, the experts were asked to determine if the current task should be further decomposed to make it actionable. For the example test case shown above, its label is No as it is actionable and does not require further task decomposition.
Prompting techniques Accuracy Statistics
Mean SD $p$ Sig.
Zero-shot .35 .000
Few-shot .58 .040
+ CoT .62 .050 .405 -
+ CoT + Tree .69 .020 .004 **
+ CoT + Draft .72 .020 .009 **
+ CoT + Tree + Draft .87 .040 .000 ***
[…]
4.1.3. Results and Findings
As summarized in Table 1, the zero-shot method yielded the lowest accuracy of only 0.35, which also implies an inherent difficulty in the task itself. We observe enhanced performance with few-shot prompting (0.58), with accuracy rising even more when using the CoT prompting paradigm (0.62). Combining CoT few-shot prompting with the task node tree levels (0.69), initial solution draft (0.72), or both (0.87), all significantly led to enhanced performance. Generating the initial solution draft introduced a trade-off between latency and accuracy, so we opted for the slightly less performant few-shot CoT with tree levels for a better user experience.
4.2. Comparative Study with and without Context Curation
To evaluate how context curation affects the quality of JumpStarter’s proposed plans and working solutions, we conducted a controlled within-subjects lab study for the following conditions:
  1. (1)
    Context dumping: context saving only.
  2. (2)
    Context filtering: context saving and context selection.
  3. (3)
    Context curation: context saving, context selection, and context elicitation.
Besides ablating the use of both context selection and context elicitation in Condition (1) and context elicitation alone in Condition (2), all features and UI were kept exactly the same. We hypothesized that action plans and working solutions developed with the full context curation method, represented by Condition (3), would be judged as higher quality than those created without it.
4.2.1. Participants and Procedure
Participant Personal goal Condition order
E1 Apply to a fellowship (1) $\rightarrow$ (2) $\rightarrow$ (3)
E2 (2) $\rightarrow$ (3) $\rightarrow$ (1)
E3 Get a driver’s license (3) $\rightarrow$ (1) $\rightarrow$ (2)
E4 (1) $\rightarrow$ (2) $\rightarrow$ (3)
E5 Organize a team event (2) $\rightarrow$ (3) $\rightarrow$ (1)
E6 (3) $\rightarrow$ (1) $\rightarrow$ (2)
Table 2. Overview of expert participants for the comparative study. Six experts were assigned one of three goals to evaluate under all three conditions, which were presented in shuffled order to avoid biasing the results.
We evaluated the effect of context curation on subtask and working solution quality for three preselected personal goals inspired by (Little, 1983): (a) Apply to a fellowship, (b) Get a driver’s license, and (c) Organize a team event. We then used a university mailing list and word of mouth to recruit participants. Each goal was assigned two “expert participants”—those who reported achieving the goal in the past six months. Overall, we recruited six expert participants for this study (average age=25.8, three female, three male). Participants were compensated \$20 per hour, with sessions lasting about 1.5 hours each.
During the study, we first introduced the expert participants to the concept of action plans and working solutions with examples. We instructed the participants that they would be using three versions of the same system for their respective goal, from start to finish. The system versions represented the three experimental conditions, which were presented in shuffled order among participants to counterbalance the learning effect (see Table 2). We demonstrated how to use each version before each participant used it. We asked the participants to generate the subtasks and working solutions exactly once. They then rated the perceived quality of the subtasks and working solutions on a seven-point Likert scale, providing a verbal explanation for their choices.
4.2.2. Results and Findings
Subtask quality Working solution quality
[…]
Overall, 46 subtasks, or plan items, were generated for Condition 1 (context dumping: context saving only), 50 for Condition 2 (context filtering: context saving and context selection), and 50 for Condition 3 (context curation: context saving, context selection and context elicitation). An equal number of corresponding working solution drafts were also produced, as we asked participants to generate a draft only once for each subtask. Our method of context curation, represented by Condition (3), performed the best for both subtask quality and working solution quality. For subtask quality, context curation ($\mu=6.12$, $\sigma=0.95$) was rated higher than the similarly performing context dumping ($\mu=5.26$, $\sigma=1.26$) and context filtering ($\mu=5.28$, $\sigma=1.25$) methods. For the quality of working solutions, context curation ($\mu=6.36$, $\sigma=0.95$) outperformed context filtering ($\mu=5.68$, $\sigma=1.01$) by a similar margin, which also outperformed context dumping ($\mu=5.04$, $\sigma=1.04$). These results and corresponding p-values are reported in Table 3.
Our results show that the context selection feature significantly enhances the quality of working solutions compared to the baseline. Unlike context saving only, which keeps all context in the context window all the time, context selection requires the LLM to explicitly choose relevant context from the available pool. During sessions under this condition, participants often remarked that the generated solutions appeared to take into account what they had input in previous subtasks. This was particularly evident in “summarizing” tasks, where the system could provide a personalized checklist for tasks like applying for scholarships and driver’s licenses, or an overall itinerary for a team event. E5 referred to the event itinerary they got as a “very useful synthesis of everything I’ve explored.” In contrast, the context saving only condition tended to produce only general tips for creating the itinerary. As another example, E1 mentioned that the email draft generated with context selection was more personalized than that created without it. E1 stated, “I like the recommendation letter request email draft it gives me, as it considers much of my background that I saved in the previous ‘update your CV’ task. I did the same thing in the previous round [context dumping] but did not feel it was as effective.”
In addition to context selection, context elicitation improves the quality of both subtasks and working solutions, outperforming both context dumping and context filtering conditions. Participants reported that the elicitation questions posed at the beginning “provided the right plan to start with.” (E4) For instance, E2 uploaded the fellowship requirements document as initially suggested by the system, later rating the generated subtasks a perfect 7/7. “It captured the requirements quite accurately…” E2 noted. “The subtasks were precise, fitting the unique aspects of the fellowship I am applying to, which requires only one recommendation letter, though typically more are needed.” Similarly, elicitation questions about which state to obtain a driver’s license in (for the “get a driver license” goal) and how many people are in the team (for the “organize a team event” goal) both resulted in subtasks that were better tailored to participants’ individual situations. In addition, preference elicitation questions such as “What type of vehicle do you intend to drive?” and “When do you prefer to hold the event? Weekday or weekend? Noon or night?” prompted participants to provide answers as personal context, ultimately resulting in more tailored working solutions that they rated highly.
5. User Study
To understand how JumpStarter might assist users in getting started on their personal goals, we conducted a within-subjects study with ten participants, comparing JumpStarter to ChatGPT. We chose ChatGPT as the baseline because it is a popular AI tool that many people are familiar with, and it allows users to feasibly add contextual information throughout the chat session.
During the study, each participant was asked to explore one personal goal they would like to achieve in the near future using each system. After finishing the exploration with each system, participants were asked to evaluate the system through questionnaires on task load, outcome satisfaction, and confidence level in taking the next steps on their goal. We also conducted semi-structured interviews to understand their experiences with each system. At the end of each study, we conducted another interview to get feedback about their overall experience and system preferences.
5.1. Hypotheses
In the study, we investigate the following hypotheses:
  1. H1
    Compared to ChatGPT, JumpStarter significantly lowers users’ task load (H1) for mental demand (H1a), temporal demand (H1b), performance (H1c), effort (H1d) and frustration (H1e).
  2. H2
    Compared to ChatGPT, JumpStarter significantly increases users’ exploration efficiency (H2) in terms of the number of plan items explored (H2a) and the number of working solutions drafted (H2b) within the given time.
  3. H3
    Compared to ChatGPT, JumpStarter significantly increases users’ satisfaction level with the quality of results (H3) in terms of plan quality (H3a) and working solution quality (H3b).
  4. H4
    Compared to ChatGPT, JumpStarter significantly increases users’ confidence in taking the next steps on their personal projects (H4).
5.2. Participants and Procedure
Participant Personal goal Goal type
P1 Start a side job Career
P2 Organize a weekly game night Life
P3 Land a job offer Career
P4 Prepare for the LSAT Academia
P5 Manage social media accounts Creativity
P6 Move to a new apartment Life
P7 Create a portfolio website Creativity
P8 Prepare to deliver a tutorial Academia
P9 Start a YouTube channel Creativity
P10 Organize a family reunion Life
Table 4. Overview of personal goals picked by participants in the user study.
We recruited ten participants (average age=23.8, six female, four male) through a university mailing list and word of mouth. Each participant reported that they are familiar or very familiar with the use of ChatGPT. Before the study, we asked each participant to pick one personal goal they wanted to achieve in the next six months. The personal goals that each participant picked can be found in Table 4.
During the session, we first introduced participants to the concept of action plans and working solutions with examples. Then, each participant was asked to create an action plan for their goal and draft working solutions for as many plan items as possible, using ChatGPT and JumpStarter. Participants were randomly assigned to a condition order (ChatGPT first and then JumpStarter, or JumpStarter first and then ChatGPT) that was counterbalanced to prevent a learning effect. Before participants started using JumpStarter, we used an example case to guide them through the system and familiarize them with the interface. Then, participants were given a maximum of 25 minutes to complete each task. They also had access to a Google Doc to note their results if needed. The entire process took approximately 1.5 hours. Participants were compensated \$20/hour.
5.3. Results and Findings
Category Factor JumpStarter ChatGPT Statistics Hypotheses
Mean SD Mean SD $p$ Sig.
Task load Mental demand 3.3 1.64 5.3 1.06 .005 ** H1a accepted
Temporal demand 2.4 .97 4.7 1.83 .012 * H1b accepted
Performance 2.7 .95 4.6 1.26 .007 ** H1c accepted
Effort 2.8 1.69 4.8 1.23 .007 ** H1d accepted
Frustration 2.1 1.60 4.5 1.90 .009 ** H1e accepted
Exploration efficiency Plan items explored 4.5 1.43 2.9 .74 .002 ** H2a accepted
Working solutions drafted 5.3 1.89 3.3 1.16 .047 * H2b accepted
Satisfaction Plan quality 5.9 .88 3.3 1.49 .004 ** H3a accepted
Working solution quality 5.8 .92 4.7 1.25 .017 * H3b accepted
Confidence 5.6 1.07 3.9 .99 .007 ** H4 accepted
[…]
We collected participants’ ratings on a 1-7 point scale through questionnaires (see Table 5 for results). Participants rated their task load, outcome satisfaction, and confidence level in taking the next steps on their goal by using JumpStarter and ChatGPT. We recorded the number of plan items explored and the working solution drafted by each participant in each task. During the interviews, we asked them follow-up questions to understand the reasons behind their scores. We applied the thematic analysis method (Braun and Clarke, 2006) to analyze the interview transcripts. We report the key findings in this section.
5.3.1. Task load
In the NASA TLX dimensions (Hart, 2006), working with JumpStarter was significantly less demanding in terms of mental demand (p=.005, Z=-2.52), temporal demand (p=.012, Z=-2.26), performance (p=.007, Z=-2.47), effort (p=.007, Z=-2.45), and frustration (p=.009, Z=-2.34).
JumpStarter reduces the mental load by providing an easier way for users to visualize the structure and track progress, unlike ChatGPT’s linear chat interface. For example, P2 commented, “ChatGPT info dumps a lot, and I have to keep the structure in my brain, whereas JumpStarter gave me a structure that I could easily follow.” Additionally, JumpStarter eases users’ mental load by posing targeted questions to help them specify their context. For instance, P10 noted, “I appreciate the questions the system (JumpStarter) asked when I felt stuck about how to iterate on drafting events [for a family reunion]. Entering the ages of the family members actually provided many better choices.” In contrast, in ChatGPT, users must devise the context themselves and frame the prompts accordingly, as P8 said, “In ChatGPT, the information load is high—I have to think very hard about what information I should provide in order to get things that work for me.”
Moreover, participants reported putting significantly more effort into the task while using ChatGPT, although this did not necessarily result in better performance. For example, feeling that the current response was too general and niche, P1 asked ChatGPT to generate questions they could answer to improve the response. ChatGPT then provided eight questions, which were, according to P1, “hard to answer as there were many and they were a bit too abstract.” However, even after P1 made lots of effort to answer all eight questions, they still felt that the responses were “a bit too general and not useful.” Additionally, P1 noted that “ChatGPT seemed to forget these eight answers soon after in my following chat with it,” leaving them feeling that their efforts were not valued as ChatGPT failed to maintain and use this context.
5.3.2. Exploration efficiency
During the study session, participants explored significantly more plan items using JumpStarter (mean=4.5, SD=1.43) than they did with ChatGPT (mean=2.9, SD=0.74), and they drafted significantly more working solutions with JumpStarter (mean=5.3, SD=1.89) compared to ChatGPT (mean=3.3, SD=1.16).
In ChatGPT, it can be easy for users to become fixated on a single plan item, whereas JumpStarter helps users maintain an overview of the entire plan. For example, P2 went through nine iterations of the survey questions for organizing the weekly game night, stating, “I got the survey questions I like in the end [after using ChatGPT], but it took me nine iterations on my own. I really hoped ChatGPT would guide me to the right place, but I had to direct myself. Because I got too involved in this survey thing, I totally forgot I needed to work on other planning items.” Similarly, P8 spent all the time digging into details like designing the visual content for tutorial slides. They reflected “[With ChatGPT,] I went into a detailed level that was not necessary very quickly, and I just forgot to do the general planning.”
Furthermore, JumpStarter’s structure enables users to jump between tasks and focus on one manageable task at a time, making it easier for them to arrive at a working solution. P4 summarized it well: “I feel JumpStarter has a more flexible structure compared to ChatGPT. I like that I can easily jump between tasks and pick the ones I care about the most. Also, I like the visualization for each task—seeing the title, description, and context provides everything I need and helps me focus on exploring that task.” P10 echoed this sentiment, commenting, “JumpStarter automatically manages and considers my drafts of previous tasks, which is good because then I can focus on the current task.”
5.3.3. Perceived quality of outcome
Plans created with JumpStarter were perceived to be of significantly higher quality (mean=5.9, SD=0.88) compared to those created with ChatGPT (mean=3.3, SD=1.49). Moreover, the working solution quality was also perceived to be significantly higher with JumpStarter (mean=5.8, SD=0.92) compared to ChatGPT (mean=4.7, SD=1.25), based on participants’ self-reported results.
JumpStarter often poses important questions to clarify users’ needs and incorporates that information into its planning. While ChatGPT tends to provide more general plans that might not align with specific user situations. For example, P9, who wanted to start their YouTube channel, shared, “Identifying the target audience was not listed while I was using ChatGPT—I had to ask about it separately. Meanwhile, in JumpStarter, I was asked at the very beginning if I already knew my target audience. I said I needed help, and then it listed identifying the target audience as the first step in my plan, which was nice.” In addition, P9 mentioned that the plan generated with ChatGPT focused heavily on video editing, but as a video editing expert, they did not require much preparation for that task. Instead, “JumpStarter asked about my familiarity with video editing at the very beginning, and I told them I was an expert. I feel that this was reflected well in the plan—like it assigned a shorter duration for the video editing task, which better suited my situation.”
5.3.4. Confidence level in taking the next steps on the goal
Participants reported significantly higher confidence in taking the next steps on their personal projects using JumpStarter (mean=5.6, SD=1.07) compared to using ChatGPT (mean=3.9, SD=.99).
Participants reported that ChatGPT helped validate their thoughts with commonsense knowledge and sometimes provided surprising or useful tips. As P7 said, “Sitting down and planning things out itself is very helpful. I used ChatGPT as a cross-reference, checking to make sure I’m on the right track—thinking similarly to other people. And sometimes answers to low-level tasks covered things I did not really know, which is good.” However, they reported that they still always felt they might miss something important while using ChatGPT, whereas with JumpStarter, they feel more secure (P3, P4, P9), as P4 commented, “I love that I can break things down further if I want, so I don’t feel like I miss anything.”
In addition, JumpStarter can provide very personalized and actionable next steps that greatly increase users’ confidence in taking action on their projects. For example, P1, with the goal of starting a side job, shared that “The schedule JumpStarter helped me generate is very personalized, and I can directly use it to take real action—before, I felt worried about launching this idea as I had very limited time, now I feel like I can really start doing it.” P10 also liked that JumpStarter provided them with a specific and personalized itinerary for organizing the family reunion—“I like that it summarizes everything I saved in the previous tasks—I can use it in the real world.”
5.3.5. Tool preference
8 out of 10 participants reported they prefer to use JumpStarter in the future, compared to ChatGPT. The main reasons given include that JumpStarter can provide more customized responses with less cognitive load. The users feel that they do not have to think hard about what information to provide (P1), are guided by the system (P2), can more easily consume the information (P3) or track the plan (P6), and can get their personal details efficiently organized, framed, and utilized (P8, P9, P10).
The other two participants (P4 and P5) mentioned that their choice depended on how familiar they were with the project they wanted to work on. If it was a topic they already had a clear understanding of, they preferred the chatbot interaction to help them figure out the details. Otherwise, they would prefer to use JumpStarter as it offers more structure. We further our discussion in terms of choices between structure and chat in Section 6.2.
5.3.6. Improvement feedback
Participants also provided insights on how to improve JumpStarter. The main feedback included “make the subtask outline and task descriptions editable” (P1, P3, P6, P7, P8, P9), “format the suggested working solution to be easier to read” (P4, P8, P9), “enable users to add or edit context whenever they want” (P3, P7), and “add a synthesis button to summarize what has been explored so far” (P1, P2). P8 also suggested embedding a search agent to collect data and ensure credibility. We discuss limitations and future work further in Section  6.3.
6. Discussion
6.1. Adpative Personal Context Curation in Complex Problem Solving
JumpStarter’s use of adaptive personal context curation in helping users explore complex goals provides a design paradigm for future interfaces. Systems that effectively respond to personal context can better collaborate with users in achieving their goals, as demonstrated in our user studies. In JumpStarter, context curation involves not only saving necessary information but also incorporates two key components: context elicitation and context selection. We found that soliciting and integrating relevant context is essential for devising personalized plans and solutions of high quality.
Users greatly appreciated the elicitation questions posted by the system. Compared to the “pull” mode in ChatGPT, where users had to identify the requirements and constraints themselves, they tend to prefer the ”push” mode in JumpStarter, where they can review the posted questions and select which ones to answer. This preference aligns with cognitive psychology literature, suggesting that recalling information is more cognitively demanding than recognizing it (Craik and Lockhart, 1972). Additionally, when users invest more effort into prompting, as observed in ChatGPT, they are more likely to be critical of the responses they receive. This exemplifies the expectation confirmation theory (Oliver, 1980), which suggests that dissatisfaction arises when results do not meet the standards anticipated based on the effort invested.
Users also appreciated the system’s assistance in selecting and previewing context before drafting a solution, giving them a good sense of transparency and control. However, a few participants expressed a desire for the system to remember and use all the context they had provided. Tools like ChatGPT are designed to maintain and utilize a memory of all user-provided context throughout the chat session (Ray, 2023). However, our observations reveal that ChatGPT did not effectively use the provided context to deliver more personalized responses, and sometimes even provided conflicting responses to the user. This presents an interesting challenge  (Wang et al., 2024)—as users progress in their problem-solving journey, their answers to some context, such as the amount of time they want to allocate each week for LSAT preparation, might change as they realize the extent of content they need to cover. In such cases, it is crucial for the system to update the user preferences in the context pool. We aim to explore ways to refine context selection and integration from an existing context pool in the future.
Adaptive personal context curation advances the vision of flexible and personalized systems that actively respond to user intent. Currently, in JumpStarter, all context curation occurs within the system—users must manually upload files or input information rather than integrating content from existing tools like email inboxes and desktop files. If more fully integrated into users’ daily workflows, JumpStarter could solicit and integrate more relevant context, providing even more actionable and useful support to users. In line with earlier explorations (Kaptelinin, 2003; Kersten and Murphy, 2006), it would be ideal if the system could perform “in-situ” context capturing, such as an add-on that integrates into users’ existing workflows, as suggested by the theory of invisible technology (Bergman and Norman, 2000). Such an add-on could directly place the email drafts it suggests into the email application as editable drafts, ready to be refined and sent. Ultimately, JumpStarter’s use of adaptive personal context curation provides a design paradigm for future systems, where human insight and computational assistance can be more effectively blended together to achieve goals.
6.2. Blending Chat and Structure for Interacting with LLMs
Since the advent of ChatGPT, chatbots have become a mainstream method for interacting with LLMs. However, our user study indicates that a structured approach is beneficial for handling more complex tasks. This is in line with findings from recent HCI work (Ma et al., 2024), which shows that structure aids in generating structured thoughts and personalized responses. Furthermore, a structured format helps users to more easily understand the information provided by GPT, avoiding the overwhelming experience of endless scrolling, as suggested in recent studies (Suh et al., 2023; Jiang et al., 2023). The subtask tree structure in our system additionally aids users in grasping the overall scope of tasks, tracking their progress, managing context, and focusing on one individual task at a time.
However, does this mean that we should always prioritize structure over chat when using LLMs for complex problem-solving? During our studies with JumpStarter, we observed that users show an interest in engaging with a chatbot at certain stages of their exploration journey. For example, when users encounter unfamiliar terms such as “research proposal” or “back-end coding” in the plan or working solutions, they prefer to consult a context-aware chatbot to understand these terms. They find the chatbot feels more natural and allows them to quickly ask follow-up questions if anything is unclear. Similarly, when users have strong pre-existing opinions about their plans, they tend to favor the chatbot for its interactivity and immediacy. This allows them to quickly receive feedback that aligns with their own thinking flow. For instance, if they have concerns about aspects such as the market size for their side job product ideas, they can immediately drill into that with the chatbot’s assistance.
In addition, our system currently uses a “form fill-in” method for users to answer the elicitation questions. Although this structured approach helps the system efficiently gather and save the context, implementing it in a chat interaction might be more natural and friendly. One everyday metaphor is when a doctor wants to learn more about your health condition, they can either give you a long form to fill in or ask in a conversational way that might make you feel more comfortable. They can follow up immediately if you find some questions unclear or if the doctor finds your answers not making sense. However, structured interaction has many benefits that chatbot interaction does not provide, as we discussed in our study findings. It is an interesting future direction to explore how we might blend structure and interaction together to complement their pros and cons and provide a better way to interact with LLMs.
6.3. Limitations and Future Work
JumpStarter primarily assists users in figuring out “how” to achieve their personal goals, prompting real-world actions by increasing simplicity, as suggested by the Fogg Behavior Model (Fogg, 2009). However, the “why” problem—motivation (such as self-regulation or emotional challenges)—is another very important dimension that JumpStarter does not address. For example, one participant shared that their confidence dropped after exploring the goal, as they realized how much they had to do to make it happen. Emotional support at this time would be very valuable. But as prior studies suggested (Bhattacharjee et al., 2024), we should be very careful about using LLMs to provide emotional support, which deserves further investigation.
JumpStarter was designed to support people with personal goals that primarily involve cognitive or knowledge work, encompassing a wide range of goals across academia, career, creativity, and life (see examples in Table 4). However, there are many other personal goals that JumpStarter does not cover, such as physical goals (e.g., losing weight), behavioral goals (e.g., overcoming shyness), and spiritual goals (e.g., coming to terms with one’s faith), as outlined in (Little, 1983). Many of these goals require motivational support, a topic explored in previous HCI works (Consolvo et al., 2009; Ekhtiar et al., 2023). Future iterations of JumpStarter could potentially expand to include these areas to accommodate a broader spectrum of personal goals.
JumpStarter utilizes GPT-4 as its core engine for providing information. While LLMs like GPT-4 are adept at synthesizing information from the Internet and can sometimes offer valid and useful references, they are also prone to generating inaccurate or hallucinated information (Achiam et al., 2023). Hence, it is crucial to integrate search agents into our system to enhance the credibility of the information provided. Also, when pursuing creative goals such as starting a YouTube channel, users require support in collecting and analyzing real-world data. Developing a search agent specifically tailored to platforms like YouTube could be a valuable direction to explore.
JumpStarter supports users from planning to initiating their projects by creating actionable prototypes. This helps users grasp the full scope of their tasks and make tangible progress. While JumpStarter does not perform tasks like sending emails, it prepares and organizes essential information for action. This initial drafting stage is crucial for overcoming the planning fallacy, where users often underestimate the complexity and time required for tasks, leading to execution difficulties (Ahmetoglu et al., 2021). Moving forward, our focus will be on assisting users throughout their project’s lifespan. Future research will explore strategies for maintaining momentum and adapting to changes during real-world execution.
JumpStarter currently requires users to manually add context to the system. As discussed in Section 6.1, adopting a richer and more flexible method for capturing context that integrates into users’ current workflows would be beneficial. Additionally, JumpStarter currently can only handle context in text form, but many context are not in text, such as YouTube channel icons or slide visual layouts. As visual understanding models like GPT-4V (Achiam et al., 2023) evolve, it would be interesting to expand JumpStarter to handle multimodal context.
7. Conclusion
In this work, we present JumpStarter, a system designed to support people in getting started with their personal goals through adaptive personal context curation. Our technical evaluation indicates that context curation plays a crucial role in generating plans and working solutions of higher quality. Our user study reveals that, compared to ChatGPT, JumpStarter enables users to explore their goals more efficiently and with less mental effort. We discuss the design implications of incorporating adaptive personal context curation into complex problem-solving. We also provide insights into how we might blend structured and conversational approaches to better support people’s interaction with LLMs.
Appendix A GPT-4 Prompts used in JumpStarter
A.1. Context Elicitation for the Goal
See Figure 8.
Figure 8. The prompt for Context Elicitation for the goal.
A.2. Context Elicitation for Drafting Working Solutions
See Figure 9.
Figure 9. The prompt for Context Elicitation for drafting working solutions.
A.3. Context Selection
A.3.1. Context Selection for Drafting
See Figure 10.
Figure 10. The prompt for Context Selection for drafting working solutions.
A.3.2. Context Selection for Forking
The prompt is as follows:
  • My user has a main purpose: {main_purpose}. My user is working on the task {task_name}: {task_description}. My user needs to break down the task into sub-tasks. Here is the current context history from the user: {context_history}. Please select the most relevant context key from the current context history that can be used to better decompose the current task into several sub-tasks for the user to get started. Do not help the user to break down the task. Please also provide explanations. Format the response like this: ¡context_key¿: ¡reasons¿. Replace the context_key with the actual key in the context history.
A.4. Subtask Suggestion
See Figure 11.
Figure 11. The prompt for Subtask Suggestion.
A.5. Subtask Detection
See Figure 12.
Figure 12. The prompt for Subtask Detection.
A.6. Fork Detection
See Figure 13.
Figure 13. The prompt for Fork Detection.
A.7. Working Solution Drafts
The prompt to generate the working draft is shown as follows:
  • My user has a main purpose: {main purpose}. Please consider the following context information from my user: {user_context}. My user needs help with the current task {current task}: {task description}
Appendix B Prompts for Technical Evaluation of Subtask Detection
B.1. Zero-shot Prompting
The prompt for zero-shot for the task of subtask detection is demonstrated below:
  • System prompt: You are a useful assistance to detect if the current task needs to be further decomposed if it is not actionable and the primary goal of the task can not be viewed as a singular, distinctive deliverable. Based on the user prompt, please output Yes if it needs to be decomposed; No otherwise meaning it is actionable and does not require task decomposition.
  • User Prompt: My user is working on the task {task title}: {task description}. My user needs to know if the current task needs to be decomposed.
B.2. Few-shot Prompting
The prompt for few-shot-only prompting is shown in Figure 14. Note that we used three in-context examples in the prompt.
Figure 14. The few-shot-only prompt for Subtask Detection.
B.3. Few-shot + CoT
We constructed the prompt in a Chain-of-Thought fashion, where GPT-4 is instructed to first generate the reasoning and then the answer. The prompt is shown in Figure 15.
B.4. Few-shot + CoT + Draft
We experimented with incorporating both CoT and the initial working solution draft into the prompt. The system prompt is shown in Figure 16. For the user prompt, before detecting subtasks. we first generated the initial working solution draft for the current task. The user prompt is shown below:
  • My user is working on the task {task title}: {task description}. The GPT response to the task is: {Draft}. My user needs to know if the current task is specific and actionable.
B.5. Few-shot + CoT + Tree + Draft
To construct the system prompt for this setting, we incorporate the tree level of each task into the prompt. The prompt is shown in Figure 17. Additionally, for the current task at hand, its tree-level information is also presented in the user prompt, as shown below:
  • My user is working on the task {task title}: {task description}. The current node level of the task is {level}. The GPT response to the task is: {Draft}. My user needs to know if the current task is specific and actionable.
Figure 15. The prompt for few-shot + CoT for Subtask Detection.
Figure 16. The prompt for few-shot + CoT + Draft for Subtask Detection.
Figure 17. The prompt for few-shot + CoT + Tree + Draft for Subtask Detection.